OpenStudy (anonymous):

Dan and Omar had a contest to see who could read more books in November. Dan read 4 1/6 books. Omar read 1 5/6 books. How many more books did Dan read than Omar? a. 2 1/3 b. 1 1/6 c. 2 1/4 d. 1 2/3

OpenStudy (anonymous):

Can u go back to the old question u posted?

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

25/6-11/6 = 14/6 = 7/3 = 2 1/3

OpenStudy (anonymous):

the answer is not in the choices but 2 1/3 is right

OpenStudy (anonymous):

a and c is the same though u sure u didn't mistype 1/4 from 1/3?

OpenStudy (anonymous):

Im telling u its @*pariis who did that not me

OpenStudy (anonymous):

hero what do you think

hero (hero):

I think you should tell me where you are stuck that way I can help you with that part.

OpenStudy (anonymous):

for its subtract

OpenStudy (anonymous):

7/6 -5/12

hero (hero):

How did you get that?

hero (hero):

If you can explain what methods you used to get that, I might be able to help you get through your confusion.

OpenStudy (anonymous):

let me do it again i put the wrong numbers

OpenStudy (anonymous):

2/12 -10/12 =2 4/12 divide 4 /12 = 2 1/3